While exploring a cave, a spelunker starts at the entrance and moves the following distances. She

goes 75.0 m north, 250 m east, 125 m at an angle 30.0° north of east, and 150 m south. Find the resultant displacement

from the cave entrance.


1
Expert's answer
2021-09-30T10:20:14-0400

Find the resultant displacement.

"A_N=75\text{ m},\\
A_E=0,\\
B_N=0,\\
B_E=250\text{ m},\\
C_N=125\sin30°,\\
C_E=125\cos30°,\\
D_N=-150 \text{ m}.\\
D_E=0.\\\space\\
R_N=A_N+B_N+C_N+D_N=-12.5\text{ m},\\
R_E=A_E+B_E+C_E+D_E=358\text{ m}.\\
R=\sqrt{R_N^2+R_E^2}=358\text{ m}.\\\space\\
\theta=\arctan\frac{|R_N|}{R_E}=2°\text{ S of E}."